# Lightweight Transformer via Unrolling of Mixed Graph Algorithms for Traffic Forecast

### v1

@@ Abstract
To forecast traffic with both spatial and temporal dimensions, we unroll a mixed-graph-based optimization algorithm into a lightweight and interpretable transformer-like neural net. Specifically, we construct two graphs: an undirected graph ${\mathcal{G}}^{u}$ capturing spatial correlations across geography, and a directed graph ${\mathcal{G}}^{d}$ capturing sequential relationships over time. We formulate a prediction problem for the future samples of signal ${\mathbf{x}}$, assuming it is “smooth” with respect to both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, where we design new $\ell_{2}$ and $\ell_{1}$-norm variational terms to quantify and promote signal smoothness (low-frequency reconstruction) on a directed graph. We construct an iterative algorithm based on alternating direction method of multipliers (ADMM), and unroll it into a feed-forward network for data-driven parameter learning. We insert graph learning modules for ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, which are akin to the self-attention mechanism in classical transformers. Experiments show that our unrolled networks achieve competitive traffic forecast performance as state-of-the-art prediction schemes, while reducing parameter counts drastically. Our code is available in https://github.com/SingularityUndefined/Unrolling-GSP-STForecast.
@@ 1 Introduction
We study the problem of forecasting city traffic with spatial and temporal dimensions, e.g., predicting car density at a set of intersections 5 minutes into the future given measurements in the past. Optimizing prediction here means exploiting both spatial correlations across geography and sequential relationships in time of how the past measurements inform future observations. When modeling the pairwise relationships as a graph, the problem is also called time-vertex signal processing (TVSP) [1].
@@
Among model-based methods, one approach to TVSP is to filter the signal along the spatial and temporal dimensions independently [2]e.g., employ graph spectral filters in the graph signal processing (GSP) literature [3, 4] along the irregular spatial dimension, and traditional Fourier filters along the regular time dimension. Another approach is to model the spatial correlations across time as a sequence of slowly time-varying graphs, and then process each spatial signal at time $t$ separately, where the graph changes over time are constrained by the Frobenius norm [5], $\ell_{1}$-norm [6], or low-rankness [7] of consecutive adjacency matrices. However, neither approach exploits spatial and temporal relations simultaneously for optimal performance.
Leveraging the prowess of modern deep learning (DL) models like transformers [8], recent data-driven efforts address TVSP by building elaborate transformer architectures to capture spatial and temporal relations for prediction [9, 10, 11]. However, these schemes are inherently uninterpretable “black boxes” requiring large training data to tune millions of parameters. They are also fragile to covariate shift [12]: statistical mismatch between training and testing data.
@@
Graph attention networks (GAT) [13] and graph transformers [14] are adaptations of the basic self-attention mechanism in transformers for graph-structured data (e.g., computing output embeddings using only input embeddings in local graph neighborhoods). However, they are still intrinsically uninterpretable and maintain large parameter sizes.
In contrast, algorithm unrolling [15] offers an alternative hybrid model-based / data-driven paradigm; by “unrolling” iterations of a model-based algorithm minimizing a well-defined optimization objective into neural layers stacked together to form a feed-forward network for data-driven parameter learning via back-propagation, the resulting network can be both mathematically interpretable and competitive in performance to state-of-the-art (SOTA) DL models. As examples, Yang et al. [16], Yu et al. [17] design optimization algorithms that unroll into transformer-like neural nets. Inspired by Yu et al. [17] and extending from Thuc et al. [18] [^1] While Thuc et al. [18] focuses on unrolling of undirected-graph-based algorithms into transformer-like neural nets for image interpolation, we extend the study to mixed graphs with both undirected and directed edges to capture spatial and temporal pairwise relations, respectively, for traffic forecast., we unroll a dual-graph-based optimization algorithm into a lightweight interpretable transformer-like neural net for the traffic forecast problem.
@@
Specifically, we propose a novel mixed-graph approach and learn undirected graph ${\mathcal{G}}^{u}$ and directed graph ${\mathcal{G}}^{d}$ from data, to capture spatial correlations across geography and sequential relationships over time, respectively. We show that our graph learning modules are akin to the basic self-attention mechanism [19], and thus our unrolled graph-based neural nets are transformers. Given the two learned graphs, we define a prediction objective for future samples in signal ${\mathbf{x}}$, assuming ${\mathbf{x}}$ is “smooth” with respect to (w.r.t.) both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ in variational terms: graph Laplacian regularizer (GLR) [20] for undirected ${\mathcal{G}}^{u}$, and newly designed directed graph Laplacian regularizer (DGLR) and directed graph total variation (DGTV) for directed ${\mathcal{G}}^{d}$. We construct a corresponding iterative optimization algorithm based on alternating direction method of multipliers (ADMM) [21] that unrolls into neural layers for parameter learning. For the first time, our proposed DGLR and DGTV variational terms quantify and promote signal smoothness on a directed graph, with intuitive low-pass filter interpretations. Experiments show that our unrolled networks achieve competitive traffic forecast performance as SOTA prediction schemes, while employing drastically fewer parameters (6.4% of transformer-based PDFormer [22]).
@@ 2 Preliminaries
Denote by ${\mathcal{G}}^{u}({\mathcal{V}},{\mathcal{E}}^{u},{\mathbf{W}}^{u})$ an undirected graph with node set ${\mathcal{V}}=\{1,\ldots,N\}$ and undirected edge set ${\mathcal{E}}^{u}$, where $(i,j)\in{\mathcal{E}}^{u}$ implies that an undirected edge exists connecting nodes $i$ and $j$ with weight $w_{i,j}^{u}=W_{i,j}^{u}$, and $(i,j)\notin{\mathcal{E}}$ implies $W_{i,j}^{u}=0$. ${\mathbf{W}}^{u}\in\mathbb{R}^{N\times N}$ is the symmetric adjacency matrix for ${\mathcal{G}}^{u}$. Denote by ${\mathbf{D}}^{u}\in\mathbb{R}^{N\times N}$ a diagonal degree matrix, where $D_{i,i}^{u}=\sum_{j}W_{j,i}^{u}$. The symmetric graph Laplacian matrix is ${\mathbf{L}}^{u}\triangleq{\mathbf{D}}^{u}-{\mathbf{W}}^{u}$ [3]. ${\mathbf{L}}^{u}$ is positive semi-definite (PSD) if all edge weights are non-negative $w^{u}_{i,j}\geq 0,\forall i,j$ [4]. The symmetric normalized graph Laplacian matrix is ${\mathbf{L}}^{u}_{n}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1/2}{\mathbf{W}}^{u}({\mathbf{D}}^{u})^{-1/2}$, while the asymmetric random-walk graph Laplacian matrix is ${\mathbf{L}}^{u}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1}{\mathbf{W}}^{u}$.
One can define a graph spectrum by eigen-decomposing the real and symmetric graph Laplacian ${\mathbf{L}}^{u}$ (or normalized graph Laplacian ${\mathbf{L}}_{n}^{u}$), where the $k$-th eigen-pair $(\lambda_{k},{\mathbf{v}}_{k})$ is interpreted as the $k$-th graph frequency and Fourier mode, respectively [3]. Given $\{{\mathbf{v}}_{k}\}$ are orthonormal vectors, one can decompose a signal ${\mathbf{x}}$ into its graph frequency components as ${\boldsymbol{\alpha}}={\mathbf{V}}^{\top}{\mathbf{x}}$, where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$, and ${\mathbf{V}}^{\top}$ is the graph Fourier transform (GFT).
The $\ell_{2}$-norm graph Laplacian regularizer (GLR) [20], ${\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=\sum_{(i,j)\in{\mathcal{E}}^{u}}w^{u}_{i,j}(x_{i}-x_{j})^{2}$, is used to regularize an ill-posed signal restoration problem, like denoising, to bias low-frequency signal reconstruction (i.e., signals consistent with similarity graph ${\mathcal{G}}^{u}$) given observation ${\mathbf{y}}$:
@@
$\mu$ is a bias-variance tradeoff parameter [23], and the low-pass filter response is $f(\lambda)=(1+\mu\lambda)^{-1}$.
@@
We define analogous notations for a directed graph, denoted by ${\mathcal{G}}^{d}({\mathcal{V}},{\mathcal{E}}^{d},{\mathbf{W}}^{d})$. $[i,j]\in{\mathcal{E}}^{d}$ implies a directed edge exists from node $i$ to $j$ with weight $w_{j,i}^{d}=W_{j,i}^{d}$. ${\mathbf{W}}^{d}\in\mathbb{R}^{N\times N}$ is the asymmetric adjacency matrix. Denote by ${\mathbf{D}}^{d}\in\mathbb{R}^{N\times N}$ a diagonal in-degree matrix, where $D_{i,i}^{d}=\sum_{j}W_{i,j}^{d}$. The directed graph Laplacian matrix is defined as ${\mathbf{L}}^{d}\triangleq{\mathbf{D}}^{d}-{\mathbf{W}}^{d}$. The directed random-walk graph Laplacian matrix is ${\mathbf{L}}^{d}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{d})^{-1}{\mathbf{W}}^{d}$.
Similar frequency notion cannot be simply defined for directed graphs via eigen-decomposition of directed graph Laplacian ${\mathbf{L}}^{d}$, because ${\mathbf{L}}^{d}$ is asymmetric, and thus the Spectral Theorem [24] does not apply. We circumvent this problem by designing new variational terms in Section 3.2.
@@ 3.2.1 $\ell_{2}$-norm Variational Term
Interpreting ${\mathbf{W}}_{r}^{d}$ as a graph shift operator (GSO) [25] on ${\mathcal{G}}^{d}$, we first define an $\ell_{2}$-norm variational term [^2] A similar directed graph variational term was defined in Li et al. [26] towards an objective for directed graph sampling. We focus instead on directed graph signal restoration. for ${\mathcal{G}}^{d}$ as the squared difference between ${\mathbf{x}}$ and its graph-shifted version ${\mathbf{W}}_{r}^{d}{\mathbf{x}}$:
@@
where symmetrized directed graph Laplacian matrix ${\mathcal{L}}_{r}^{d}\triangleq({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is symmetric and PSD. We call ${\mathbf{x}}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ the DGLR, which computes the sum of squared differences between each child $j\in\bar{{\mathcal{S}}}$ and its parents $i$ for $[i,j]\in{\mathcal{E}}^{d}$. Note that DGLR computes to zero for the constant vector ${\mathbf{1}}$:
@@
We interpret eigen-pairs $\{(\xi_{k},{\mathbf{u}}_{k})\}$ of ${\mathcal{L}}^{d}_{r}$ as graph frequencies and graph Fourier modes for directed graph ${\mathcal{G}}^{d}$, respectively, similar to eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of graph Laplacian ${\mathbf{L}}^{u}$ for an undirected graph. Importantly, we prove that our frequency definition using DGLR defaults to pure sinusoids as frequencies in the unweighted directed line graph case. Thus, our algorithm minimizing DGLR includes TVSP schemes like Ramakrishna et al. [2] that analyze signals along the time dimension using classical Fourier filters as special cases.
@@
See Appendix A for a formal proof.
@@
Remark: Using our symmetrized ${\mathcal{L}}^{d}_{r}$ does not mean directionality has been lost. For the 3-node DAG in Fig. 2(c), the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{3}-\frac{1}{2}(x_{1}+x_{2}))^{2}$ and computes to 00 for ${\mathbf{x}}=[2~{}~{}0~{}~{}1]^{\top}$. On the other hand, for the DAG in Fig. 2(d) with opposite directions, the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{1}-x_{3})^{2}+(x_{2}-x_{3})^{2}$ and computes to 00 only if ${\mathbf{x}}$ is a constant vector.

### v2

# Lightweight and Interpretable Transformer via Mixed Graph Algorithm Unrolling for Traffic Forecast

@@ Abstract
Unlike conventional “black-box” transformers with classical self-attention mechanism, we build a lightweight and interpretable transformer-like neural net by unrolling a mixed-graph-based optimization algorithm to forecast traffic with spatial and temporal dimensions. We construct two graphs: an undirected graph ${\mathcal{G}}^{u}$ capturing spatial correlations across geography, and a directed graph ${\mathcal{G}}^{d}$ capturing sequential relationships over time. We predict future samples of signal $\mathbf{x}$}, assuming it is “smooth” with respect to both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, where we design new $\ell_{2}$ and $\ell_{1}$-norm variational terms to quantify and promote signal smoothness (low-frequency reconstruction) on a directed graph. We design an iterative algorithm based on alternating direction method of multipliers (ADMM), and unroll it into a feed-forward network for data-driven parameter learning. We insert graph learning modules for ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ that play the role of self-attention. Experiments show that our unrolled networks achieve competitive traffic forecast performance as state-of-the-art prediction schemes, while reducing parameter counts drastically. Our code is available in https://github.com/SingularityUndefined/Unrolling-GSP-STForecast.
@@ 1 Introduction
Transformer, based on the classical self-attention mechanism (Vaswani et al., 2017), is now a dominant deep learning (DL) architecture that has achieved state-of-the-art (SOTA) performance across multiple fields (Woźniak et al., 2020; Zamir et al., 2022). However, like other “off-the-shelf” DL models, transformer has a massive number of parameters, and its internal structure is not easily interpretable. Parameter reduction is a major industrial concern for practical computation and/or memory-constrained environments [^1] See a brief review of lightweight learning models for different applications in Appendix A.1.. Algorithm unrolling (Monga et al., 2021) offers an alternative hybrid model-based / data-driven paradigm; by “unrolling” iterations of a model-based algorithm minimizing a well-defined optimization objective into neural layers stacked together to form a feed-forward network for data-driven parameter learning, the resulting network can be both mathematically interpretable [^2] Common in algorithm unrolling (Monga et al., 2021), “interpretability” here means that each neural layer corresponds to an iteration of an optimization algorithm minimizing a mathematically-defined objective. and competitive in performance. Notably, Yu et al. (2023) designs an algorithm minimizing a sparse rate reduction (SRR) objective that unrolls into a transformer-like neural net for image classification. Thuc et al. (2024) designs graph-based algorithms minimizing graph smoothness priors that unroll into transformer-like neural nets for image interpolation while reducing parameters. However, only a positive undirected graph model was used in Thuc et al. (2024) to capture simple pairwise correlations between neighboring pixels in a static image.
Spatial-temporal data (e.g., traffic and weather) contains more complex node-to-node relationships: geographically near stations exhibit spatial correlations, while past observations influences future data, resulting in sequential relationships. In this paper, we study the design of transformers via graph algorithm unrolling for spatial-temporal data, using traffic forecast as a concrete example application.
@@
Among model-based methods, one approach to traffic forecast is to filter the signal along the spatial and temporal dimensions independently (Ramakrishna et al., 2020)e.g., employ graph spectral filters in the graph signal processing (GSP) literature (Ortega et al., 2018; Cheung et al., 2018) along the irregular spatial dimension, and traditional Fourier filters along the regular time dimension. Another approach is to model the spatial correlations across time as a sequence of slowly time-varying graphs, and then filter each spatial signal at time $t$ separately, where the changes in consecutive graph adjacency matrices in time are constrained by the Frobenius norm (Kalofolias et al., 2017), $\ell_{1}$-norm (Yamada et al., 2019), or low-rankness (Bagheri et al., 2024). However, neither approach exploits spatial and temporal relationships simultaneously for optimal performance.
@@
Recent DL-based efforts address traffic forecast by building elaborate transformer architectures to capture spatial and temporal relations (Feng & Tassiulas, 2022; Gao et al., 2022; Jin et al., 2024) [^3] A detailed review of DL models on traffic forecast is included in Appendix A.2.. Like transformers used in other fields (Zamir et al., 2022), these are also black boxes with large parameter counts. Graph attention networks (GAT) (Velickovic et al., 2018) and graph transformers (Dwivedi & Bresson, 2021) are adaptations of the self-attention mechanism in transformers for graph-structured data (e.g., computing output embeddings using only input embeddings in local graph neighborhoods). However, they are still uninterpretable and maintain large parameter sizes.
@@
In contrast, we extend Thuc et al. (2024) to build lightweight “white-box” transformers that are fully interpretable via mixed-graph algorithm unrolling for traffic forecast. Specifically, we learn two graphs from data: i) undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations across geography, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships over time. We show that our graph learning modules play the role of self-attention (Bahdanau et al., 2014), and thus our unrolled graph-based neural nets are transformers. Given the two learned graphs, we define a prediction objective for future samples in signal $\mathbf{x}$}, assuming $\mathbf{x}$} is “smooth” with respect to (w.r.t.) both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ in variational terms: graph Laplacian regularizer (GLR) (Pang & Cheung, 2017) for undirected ${\mathcal{G}}^{u}$, and newly designed directed graph Laplacian regularizer (DGLR) and directed graph total variation (DGTV) for directed ${\mathcal{G}}^{d}$. We design a corresponding linear-time optimization algorithm based on alternating direction method of multipliers (ADMM) (Boyd et al., 2011) that unrolls into neural layers for parameter learning. For the first time, our DGLR and DGTV variational terms quantify and promote signal smoothness on a directed graph, with intuitive low-pass filter interpretations. Experiments show that our networks achieve competitive traffic forecast performance as SOTA prediction schemes, while employing drastically fewer parameters (6.4% of transformer-based PDFormer (Jiang et al., 2023)).
@@ 2 Preliminaries
Denote by ${\mathcal{G}}^{u}({\mathcal{V}},{\mathcal{E}}^{u},{\mathbf{W}}^{u})$ an undirected graph with node set ${\mathcal{V}}=\{1,\ldots,N\}$ and undirected edge set ${\mathcal{E}}^{u}$, where $(i,j)\in{\mathcal{E}}^{u}$ implies that an undirected edge exists connecting nodes $i$ and $j$ with weight $w_{i,j}^{u}=W_{i,j}^{u}$, and $(i,j)\notin{\mathcal{E}}$ implies $W_{i,j}^{u}=0$. ${\mathbf{W}}^{u}\in\mathbb{R}^{N\times N}$ is the symmetric adjacency matrix for ${\mathcal{G}}^{u}$. Denote by ${\mathbf{D}}^{u}\in\mathbb{R}^{N\times N}$ a diagonal degree matrix, where $D_{i,i}^{u}=\sum_{j}W_{j,i}^{u}$. The symmetric graph Laplacian matrix is ${\mathbf{L}}^{u}\triangleq{\mathbf{D}}^{u}-{\mathbf{W}}^{u}$ (Ortega et al., 2018). ${\mathbf{L}}^{u}$ is positive semi-definite (PSD) if all edge weights are non-negative $w^{u}_{i,j}\geq 0,\forall i,j$ (Cheung et al., 2018). The symmetric normalized graph Laplacian matrix is ${\mathbf{L}}^{u}_{n}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1/2}{\mathbf{W}}^{u}({\mathbf{D}}^{u})^{-1/2}$, while the asymmetric random-walk graph Laplacian matrix is ${\mathbf{L}}^{u}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1}{\mathbf{W}}^{u}$.
One can define a graph spectrum by eigen-decomposing the real and symmetric graph Laplacian ${\mathbf{L}}^{u}$ (or normalized graph Laplacian ${\mathbf{L}}_{n}^{u}$), where the $k$-th eigen-pair $(\lambda_{k},{\mathbf{v}}_{k})$ is interpreted as the $k$-th graph frequency and Fourier mode, respectively (Ortega et al., 2018). Given $\{{\mathbf{v}}_{k}\}$ are orthonormal vectors, one can decompose a signal $\mathbf{x}$} into its graph frequency components as ${\bm{\alpha}}={\mathbf{V}}^{\top}{\mathbf{x}}$, where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$, and $\mathbf{V}$$^{\top}$ is the graph Fourier transform (GFT).
The $\ell_{2}$-norm graph Laplacian regularizer (GLR) (Pang & Cheung, 2017), ${\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=\sum_{(i,j)\in{\mathcal{E}}^{u}}w^{u}_{i,j}(x_{i}-x_{j})^{2}$, is used to regularize an ill-posed signal restoration problem, like denoising, to bias low-frequency signal reconstruction (i.e., signals consistent with similarity graph ${\mathcal{G}}^{u}$) given observation $\mathbf{y}$}:
@@
$\mu$ is a bias-variance tradeoff parameter, and the low-pass filter response is $f(\lambda)=(1+\mu\lambda)^{-1}$.
@@
We define analogous notations for a directed graph, denoted by ${\mathcal{G}}^{d}({\mathcal{V}},{\mathcal{E}}^{d},{\mathbf{W}}^{d})$. $i,j]\in{\mathcal{E}}^{d}$ implies a directed edge exists from node $i$ to $j$ with weight $w_{j,i}^{d}=W_{j,i}^{d}$. ${\mathbf{W}}^{d}\in\mathbb{R}^{N\times N}$ is the asymmetric adjacency matrix. Denote by ${\mathbf{D}}^{d}\in\mathbb{R}^{N\times N}$ a diagonal in-degree matrix, where $D_{i,i}^{d}=\sum_{j}W_{i,j}^{d}$. The directed graph Laplacian matrix is defined as ${\mathbf{L}}^{d}\triangleq{\mathbf{D}}^{d}-{\mathbf{W}}^{d}$. The directed random-walk graph Laplacian matrix is ${\mathbf{L}}^{d}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{d})^{-1}{\mathbf{W}}^{d}$.
Similar frequency notion cannot be simply defined for directed graphs via eigen-decomposition of directed graph Laplacian ${\mathbf{L}}^{d}$, because ${\mathbf{L}}^{d}$ is asymmetric, and thus the Spectral Theorem (Hawkins, 1975) does not apply. We circumvent this problem by designing new variational terms in Section 3.2.
@@ 3.2.1 $\ell_{2}$-norm Variational Term
Interpreting ${\mathbf{W}}_{r}^{d}$ as a graph shift operator (GSO) (Chen et al., 2015) on ${\mathcal{G}}^{d}$, we first define an $\ell_{2}$-norm variational term [^4] A similar directed graph variational term was defined in Li et al. (2023) towards an objective for directed graph sampling. We focus instead on directed graph signal restoration. for ${\mathcal{G}}^{d}$ as the squared difference between $\mathbf{x}$} and its graph-shifted version ${\mathbf{W}}_{r}^{d}{\mathbf{x}}$:
@@
where symmetrized directed graph Laplacian matrix ${\mathcal{L}}_{r}^{d}\triangleq({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is symmetric and PSD. We call ${\mathbf{x}}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ the DGLR, which computes the sum of squared differences between each child $j\in\bar{{\mathcal{S}}}$ and its parents $i$ for $i,j]\in{\mathcal{E}}^{d}$. Note that DGLR computes to zero for the constant vector $\mathbf{1}$}:
@@
We interpret eigen-pairs $\{(\xi_{k},{\mathbf{u}}_{k})\}$ of ${\mathcal{L}}^{d}_{r}$ as graph frequencies and graph Fourier modes for directed graph ${\mathcal{G}}^{d}$, respectively, similar to eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of graph Laplacian ${\mathbf{L}}^{u}$ for an undirected graph. Importantly, we prove that our frequency definition using DGLR defaults to pure sinusoids as frequencies in the unweighted directed line graph case. Thus, our algorithm minimizing DGLR includes GSP schemes like Ramakrishna et al. (2020) that analyze signals along the time dimension using classical Fourier filters as special cases.
@@
See Appendix B for a formal proof.
@@
Remark: Using our symmetrized ${\mathcal{L}}^{d}_{r}$ does not mean directionality has been lost. For the 3-node DAG in Fig. 2(c), the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{3}-\frac{1}{2}(x_{1}+x_{2}))^{2}$ and computes to 0 for ${\mathbf{x}}=[2~~0~~1]^{\top}$. On the other hand, for the DAG in Fig. 2(d) with opposite directions, the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{1}-x_{3})^{2}+(x_{2}-x_{3})^{2}$ and computes to 0 only if $\mathbf{x}$} is a constant vector.